# Inductive Graph Representation Learning with Quantum Graph Neural Networks

### v1

@@ 6.2 Comparison Between Classical and Quantum Models
We performed numerical simulations to compare classical GNNs with our proposed QGNNs using two cases based on molecular size and sample count, evaluated by Loss and R² Score. These cases include
  • Case 1: Samples of $30$ molecules with up to 9 atoms,
  • Case 2: Samples of $20$ molecules with up to 18 atoms.
@@
Training uses 70% of the dataset, with 30% reserved for testing. All experiments are conducted over 300 epochs and the results are presented in Fig. 5 for each case, showing the training and testing results of the GNN and QGNN frameworks. Table 1 summarizes the final results obtained.
For Case 1, the GNN achieves superior training performance, while the QGNN demonstrates slightly better generalization on the test set. In Case 2, which includes molecules with up to twice as many atoms as Case 1, the classical model again outperforms during training, but the quantum model shows marginally better generalization. It is important to note that, despite the significant increase in the number of atoms from Case 1 to Case 2, the quantum model remains constant in depth and width (same number of input qubits), keeping the number of tunable parameters unchanged. This highlights its potential robustness in handling larger graphs without requiring design modifications.
However, while the classical GNN consistently outperforms the QGNN during training, the quantum approach exhibits faster or equal initial learning, particularly within the first 50 to 70 epochs. Furthermore, as molecular complexity increases, the quantum model maintains similar ability to generalize to unseen data despite declining training performance. This decrease, observed in both classical and quantum models, is likely due to the fixed tunable parameters for Cases 1 and 2.
The effectiveness of our framework in both training and testing highlights the potential of our QGNN to handle molecules of varying atom sizes, despite its minimal design. In addition, by benchmarking the framework on non-trivial datasets as QM9 and utilizing non-trivial Hamiltonian matrices (see Ref. [57]), we move beyond the simple datasets typically used to evaluate QML models in the literature. This motivates further exploration of such quantum convolutional architectures beyond classically simulable problems [27].
Case Model # Params. Train R² Train Loss Test R² Test Loss
Case 1 GNN 209 0.996 0.003 0.732 0.300
QGNN 113 0.944 0.040 0.771 0.293
Case 2 GNN 209 0.985 0.009 0.723 0.127
QGNN 113 0.848 0.091 0.781 0.100
@@
Table 1: Performance comparison of GNN and QGNN across Cases 1 and 2. The table highlights the number of tunable parameters, train and test R² scores, and losses for each model, emphasizing their respective results.
@@
(c) R² score curves for Case 2
(d) Loss curves for Case 2
@@
Figure 5: Comparison of training and testing performance for GNN (classical) and QGNN (quantum) models on molecules from the QM9 dataset, with all results obtained over 300 epochs. The markers are accompanied by dashed lines in their corresponding colors: circular and diamond markers distinguish the GNN’s train and test results, respectively, while square and triangular markers represent the QGNN’s train and test results. (a) and (b) present the R² score and Loss respectively for Case 1, and (c) and (d) present the R² score and Loss for Case 2.
As is clearly shown in Fig. 5, the QGNN loss for both train and test set consistently plateaus during training across all examined cases. Low model expressibility or the presence of suboptimal local minima in the loss landscape may be contributing factors that hinder the trainability of the quantum model [58, 59]. Another potential explanation for this behavior is the presence of barren plateaus, a well-documented issue in QML [33, 60]. Recent works on QCNNs [34] though indicate that barren plateaus are absent under certain conditions  [33]. Since the QGCN circuitry is inspired by QCNNs, adopting similar strategies can advance QGNNs into a reliable and scalable framework for quantum graph learning.
Next, we investigate whether the previously discussed techniques can also effectively identify the absence of barren plateaus in graph machine learning problems.
6.3 Scalability and absence of barren plateaus in QGNNs
@@
Barren plateaus [28, 29, 30, 31, 32] pose a significant challenge to the trainability of QML models [60, 61, 62, 63, 64], as they lead to an exponential decay in the gradients of the loss function. The absence of barren plateaus implies that minimizing the loss function landscape does not require exponentially high precision, allowing expectation values to be measured with polynomial precision relative to the system size. In this section, we numerically confirm the scalability of the proposed QGNN framework, as barren plateaus are absent with increasing number of qubits. This ensures continued trainability as the model scales. As previously discussed, the number of qubits in the QGCN circuit scales linearly with node feature dimensions and remains invariant to node count. Thus, changing the number of qubits does not alter the circuit.
To infer the presence of barren plateaus, we compute the loss function partial derivative $\operatorname{Var}[\partial_{\theta_{\mu}}\mathcal{L}]$-being $\theta_{\mu}$ a variational parameter-by sampling different points in the loss landscape. According to Chebyshev’s inequality (9), the probability that the partial derivative of the cost function deviates from its mean value (which is zero) is constrained as follow [32]
@@
Inspired by Ref. [34], we study the gradient scaling of our QGCN architecture under the same conditions. Specifically, we use the convolutional layer design consisting of independent (uncorrelated) local 2-designs $W$, as shown in Fig. 4. We also construct a correlated convolutional layer where all $W$’s within each layer share the same parameters. We note this particular Ansatz significantly reduces the number of variational parameters, which may impact the model’s expressibility. Finally, the loss function is defined by local measurements, ensuring its local nature—a characteristic inherent to quantum convolutional models due to pooling layers.
Figure 6: Average variance of the loss function partial derivatives $\overline{\operatorname{Var}}\left[\partial_{\theta_{\mu}}\mathcal{L}\right]$ of the QGCN circuit. The correlated QGNN (corr) uses the same parameters for all $W$ cells in a convolution layer, while in the uncorrelated (uncorr) all cells are independent. The variance $\operatorname{Var}\left[\partial_{\theta_{\mu}}\mathcal{L}\right]$ for each variational parameter $\theta_{\mu}$ was calculated over 200 points in the landscape.
@@
The numerical results for the variance of the loss gradients are shown in Fig. 6. The plot displays the average variance $\overline{\operatorname{Var}}\left[\partial_{\theta_{\mu}}\mathcal{L}\right]=\sum^{N}_{\mu}\operatorname{Var}\left[\partial_{\theta_{\mu}}\mathcal{L}\right]/N$ across all $N$ variational parameters for each QGCN circuit (correlated and uncorrelated). Unlike previous results, we use synthetic data as input node features to allow us to easily increase their dimensionality and, consequently, the number of qubits. By sampling 200 random points on the loss landscape, we compute the average variance for each parameter $\theta_{\mu}$, treating each number of qubit case separately. Remarkably, both the correlated and uncorrelated Ansätze suggest sub-exponential scaling of the variance, confirming the absence of barren plateaus in both models. In particular, the correlated Ansatz achieves higher variances that remain nearly constant as the number of qubits increases. This suggests that introducing correlation in the parameters can significantly enhance the trainability of the QGNN model at scale, as large gradients are essential for efficient training.
Importantly, in benchmarking our QGNN, we employ a single QGCN circuit, so the absence of barren plateaus in the model also implies its absence in the QGNN framework. Here we conclude our analysis.
7 Conclusions
@@
We presented a novel Quantum Graph Neural Network framework that efficiently generates node embeddings by combining inductive representation learning with quantum circuit design. Specifically, our QGNN integrates sampling and aggregation techniques from GraphSAGE, using QGCNs as aggregators. The QGCN circuit is inspired by previous QCNN designs [24], featuring convolutional and pooling layers, while maintaining high flexibility to accommodate diverse architectures. By defining the convolutional cell, the circuit is uniquely determined, enabling the construction of tailored quantum circuits for specific problems or symmetry-preserving architectures.
@@
Benchmarking on the QM9 dataset demonstrates that the framework achieves similarly good results to the classical framework, both regarding trainability and generalization, highlighting its capability to handle non-trivial datasets. While GNNs consistently outperform during training, the QGNN achieves comparable or slightly superior generalization, particularly as molecular complexity increases. Notably, the quantum circuit architecture remains unchanged, with tunable parameters staying constant regardless of the number of nodes in a graph, showcasing good generalization and ability to handle larger graphs without requiring design modifications. Additionally, the quantum approach shows faster initial learning, indicating potential for rapid convergence.
@@ Acknowledgments
Appendix A Appendix: Pseudocode for (Q)GNN algorithm on QM9 Dataset
In this appendix, we summarize the algorithm used to benchmark the performance of classical GNNs and quantum-enhanced QGNNs on the QM9 dataset. It systematically processes molecular graphs through atom-level and molecule-level operations, capturing both local and global structural information to predict molecular properties. The pseudocode 1 illustrates, at a high level, how each molecule is processed, focusing on both molecular properties (such as the target energy $E_{LUMO}$) and atom-level features. For each atom, an 11-dimensional feature vector is provided in PyTorch Geometric [25].
In our analysis, we employ a subset of atomic features, specifically retaining indices $0,1,2,3,4,5,10$. These indices represent the atomic number, chirality indicator, atom degree, formal charge, radical electrons, hybridization state, and scaled atomic mass. In contrast, indices $[6,7,8,9]$ encode properties such as aromaticity, explicit hydrogen count, ring system membership, and valence. These features were excluded because their values are zero across the selected samples, providing no additional information.
Algorithm 1 (Q)GNN on QM9 dataset
0:  Batch of graphs $G$ with adjacency matrices $A$, node features $F$, targets $T$, model $\mathcal{M}$
0:  $true\_targets$, $predicted\_targets$, and $avg\_loss$
1:  Initialize $avg\_loss\leftarrow 0.0$
2:  Initialize $predicted\_targets\leftarrow[\,]$, $true\_targets\leftarrow[\,]$ Process each molecule in the batch
3:  for each molecule $(A,F,T)$ in batch do
4:     Initialize $previous\_output\leftarrow\mathbf{0}$
5:     Step1: Atom-level Aggregation and Concatenation
6:     for each node $v$ in $G$ do
7:        Substep 1.1: Gathering Neighboring Nodes
8:        $connected\_features\leftarrow F[A[v]]$ (if $A[v]$ is nonzero)
9:        Substep 1.2: Concatenation
10:        $combined\_features\leftarrow\text{concat}(connected\_features,previous\_output)$
11:        Substep 1.3: Atom-level Aggregation
12:        $output\leftarrow\mathcal{M}(combined\_features)$
13:        $previous\_output\leftarrow output$
14:        Append $output$ to $atom\_outputs$
15:     end for
16:     Step2: Molecule-level Aggregation and Loss
17:     $molecular\_output\leftarrow\text{mean}(\text{$atom\_outputs$})$
18:     Append $molecular\_output$ to $predicted\_targets$
19:     Append $T$ to $true\_targets$
[…]
21:  end for
22:  return  $true\_targets$, $predicted\_targets$, $avg\_loss$
The algorithm starts with Step 1: Atom-level Aggregation and Concatenation, comprising three substeps. First, in Substep 1.1: Gathering Neighboring Nodes, each node $v$ identifies its neighbors based on the graph’s adjacency matrix, gathering their corresponding features. Next, in Substep 1.2: Concatenation, these neighboring features are combined with the node’s previous output. The concatenation is done before the aggregation step, in contrast to the vanilla GNN described earlier, to ensure the tensor size remains consistent throughout the algorithm. For example, in the first iteration, the previous output is an empty vector, making this step functionally equivalent to concatenating after aggregation in the atom-level substeps. Finally, in Substep 1.3: Atom-level Aggregation, the concatenated features are processed by the tailored GCN to produce an output for each node. This output captures both the node’s intrinsic features and information from its neighboring atoms.
Once atom-level operations are complete, the process moves to Step 2: Molecule-level Aggregation and Loss Calculation. In this step, the outputs from all nodes within a molecule are aggregated using mean pooling to produce a single molecular-level output. This molecular representation is compared with the target property using a loss function, and the computed loss contributes to adjusting the model parameters. The molecular-level aggregation and loss calculation steps enable the model to capture both local and global information about the molecular structure, enhancing the tailored GCN’s performance in the regression task.
In summary, the algorithm processes the graph systematically, combining atom- and molecule-level operations to capture both structural and feature-based molecular information. The algorithm described for the classical implementation also applies in the quantum context, with the classical model replaced by a quantum circuit that processes graph data through quantum convolutional and quantum pooling layers. This modification preserves the algorithm’s overall structure while leveraging quantum operations to improve efficiency and expressive power.

### v2

@@ VI.2 Comparison Between GNNs and QGNNs
We performed numerical simulations to compare classical GNNs with our proposed QGNNs using two cases based on molecular size and sample count, evaluated by Loss and R² Score. These cases are:
  • Case 1: a sample of 30 molecules with up to 9 atoms.
  • Case 2: a sample of 30 molecules with up to 18 atoms.
@@
Each case is sampled randomly and independently, with 70% of the sample dataset used for training and 30% for testing. All experiments are conducted over 300 epochs and the results are presented in Figs. 6 and 7 for each case, showing the training and testing results of the GNN and QGNN frameworks. Table 2 reports the best test-set results, which we use to assess generalization performance. The corresponding training loss and score are also included.
FW test R² test Loss train R² train Loss
Case 1 GNN 0.7501 0.1178 0.9526 0.0530
QGNN 0.8912 0.0518 0.8525 0.1644
Case 2 GNN 0.6499 0.1951 0.8644 0.0730
QGNN 0.8801 0.0693 0.7917 0.1157
@@
Table 2: Performance comparison of GNN and QGNN frameworks (FWs) across Cases 1 and 2. The table highlights the train and test R² scores, and losses for each model and case.
For Case 1, shown in Fig. 6, the classical GNN achieves superior training performance, whereas the QGNN exhibits improved generalization on the test set. In Case 2, Fig. 7, which includes molecules with up to twice as many atoms as in Case 1, the classical model again outperforms during training, while the quantum model maintains stronger generalization. Notably, despite the substantial increase in molecular size between the two cases, the number of input qubits, and consequently the width of the quantum model, remains fixed. Moreover, within each case, the same quantum circuit is applied to all graphs, resulting in an unchanged circuit depth and consequently number of trainable parameters per case. This holds even though Case 2 exhibits greater variability in graph size, with molecules ranging from 7 to 18 atoms, suggesting a degree of robustness in handling larger graphs without requiring architectural modifications to the circuit.
Although the classical GNN consistently achieves lower training loss, the quantum model exhibits faster initial learning, particularly within the first 100 epochs in Case 1 and the first 50 epochs in Case 2. As molecular complexity increases from Case 1 to Case 2, the quantum model retains a strong ability to generalize to unseen data despite a decrease in training performance. This decrease, observed in both classical and quantum models, is likely attributable to the use of a single (Q)GCN for all aggregation steps when graphs of increased size variability are present in the dataset.
@@
Figure 6: Comparison of training and testing performance for GNN (classical) and QGNN (quantum) models on molecules from the QM9 dataset, with all results obtained over 300 epochs. The markers are accompanied by dashed lines in their corresponding colors: circular and diamond markers distinguish the GNN’s train and test results, respectively, while square and triangular markers represent the QGNN’s train and test results. (a) and (b) present the R² score and Loss respectively for Case 1.
Despite the minimal design, these results highlight the potential of the proposed QGNN to process molecules of varying sizes and to offer advantages in generalization to unseen molecular structures. By benchmarking on a non-trivial dataset such as QM9, this study motivates further investigation of quantum convolutional graph architectures. Importantly, as is clearly shown in Figs. 6 and  7, the QGNN losses for both train and test set consistently plateaus during training across the two examined cases. Low model expressibility or the presence of suboptimal local minima in the loss landscape may be contributing factors that hinder the trainability of the quantum model [anschuetz_2021_critical, bittel_2021_training].
VI.3 Multiple quantum aggregators
To address the observed limitations in trainability, we now investigate the use of multiple quantum aggregators (or QGCN circuits) in our QGNNs framework. To this end, in ad-
(a) R² score curves for Case 2
(b) Loss curves for Case 2
Figure 7: Comparison of training and testing performance for GNN (classical) and QGNN (quantum) models on molecules from the QM9 dataset, with all results obtained over 300 epochs. The markers are accompanied by dashed lines in their corresponding colors: circular and diamond markers distinguish the GNN’s train and test results, respectively, while square and triangular markers represent the QGNN’s train and test results. (a) and (b) present the R² score and Loss respectively for Case 2.
dition to Cases 1 and 2, we consider also the cases with up to 16 and 25 atoms to cover a broader range of graph sizes. For each case, we generate 40 independent random samples. Here, rather than using a single quantum circuit as the aggregator (see the schematic representation 1(b)), we assign a distinct shallow quantum aggregator per hop, such that the number of circuits matches the maximum number of hops in each sample. For instance, in the case of graphs with up to 9 atoms, we employ at maximum 9 distinct quantum circuits. Each aggregator consists of a single convolution–pooling layer, i.e., $r=[1]$, resulting in a larger set of tunable parameters and increased model expressibility compared to using a single shared circuit. Particularly, employing shallow circuits, which are easier to optimize, can enhance both the stability and efficiency of training in quantum graph machine learning tasks. This claim is supported by the results presented in Fig. 8, obtained using an initial learning rate of $0.03$. These results indicate that the QGNN benefits substantially from this multi-QGCN design. Focusing exclusively on the training loss and score, metrics that previously underperformed, as shown in Table 2 and Figs. 67, the framework now achieves lower training loss and higher training scores across all cases, including those with molecules of up to 25 atoms.
Another approach to mitigate the observed loss stagnation and trainability issues observed in Figs. 67 is to increase the number of qubits used to encode the classical node features, thereby enhancing the expressibility of the quantum model. However, this strategy may introduce additional challenges, most notably the emergence of barren plateaus, a well-documented issue in quantum machine learning that can further hinder trainability [cerezo_nature_2021_costfunc, larocca_2024_review]. Nevertheless, recent studies on QCNNs indicate that barren plateaus can be avoided under certain architectural conditions [pesah_physrevX_2021, cerezo_provable_absence_2024]. Since the QGCN circuitry is inspired by QCNNs, adopting similar design principles may advance QGNNs toward a more reliable and scalable framework for quantum graph learning.
Next, we investigate whether the previously discussed strategies can effectively mitigate or prevent barren plateaus in our proposed QGNN framework.
VI.4 Scalability and absence of barren plateaus in QGNNs
@@
Barren plateaus [anschuetz_2022_beyond, cerezo_2020_impact, larocca_2024_review, mcclean_2018_barren] pose a significant challenge to the trainability of QML models [ragone_2023_lie, holmes_2021_connecting, bermejo_2024_improving], as they lead to an exponential decay in the gradients of the loss function. The absence of barren plateaus implies that minimizing the loss function landscape does not require exponentially high precision, allowing expectation values to be measured with polynomial precision relative to the system size. Here, we numerically confirm the scalability of the proposed QGNN framework, as barren plateaus are absent with increasing number of qubits. This ensures continued trainability as the model scales. To assess the presence of barren plateaus, we evaluate the variance of the loss function gradient, $\operatorname{Var}[\partial_{\theta_{\mu}}\mathcal{L}]$, where $\theta_{\mu}$ denotes a variational parameter, by sampling multiple points across the loss landscape. According to Ref. [arrasmith_2021_equivalence], we have the following Chebyshev’s inequality
@@
Inspired by Ref. [pesah_physrevX_2021], we study the gradient scaling of our QGCN architecture under the same conditions. Specifically, as shown in the schematic representation of the circuit. 4, we consider two convolutional layer designs: an uncorrelated construction composed of independent local 2-designs $W$, and a correlated variant in which all $W$ operators within a layer share the same parameters. Finally, the loss function is defined by local measurements, ensuring its local nature due to the pooling structure of quantum convolutional models.
(a) R² scores
(b) Losses
Figure 8: Training results of the multi-circuit approach for two graph neural network architectures, namely GNN and QGNN. Subfigure (a) reports the loss values, while subfigure (b) shows the $R^{2}$ scores for molecules of varying sizes: from $\leq 9$ to $\leq 25$ atoms. We train the models for 300300 epochs and report the best results obtained.
@@
The numerical results for the variance of the loss gradients are shown in Fig. 9. The plot displays the average variance $\overline{\operatorname{Var}}\left[\partial_{\theta_{\mu}}\mathcal{L}\right]=\sum^{N}_{\mu}\operatorname{Var}\left[\partial_{\theta_{\mu}}\mathcal{L}\right]/N$ across all $N$ variational parameters for each QGCN architecture (correlated and uncorrelated). Unlike previous results, we use synthetic data as input node features to allow us to easily increase their dimensionality and, consequently, the number of qubits. By sampling 200 random points on the loss landscape, we compute the average variance for each parameter $\theta_{\mu}$, treating each number of qubit case separately. Remarkably, both the correlated and uncorrelated Ansätze suggest sub-exponential scaling of the variance, confirming the absence of barren plateaus in both models. In particular, the correlated Ansatz achieves higher variances that remain nearly constant as the number of qubits increases. This suggests that introducing correlation in the parameters can enhance the trainability of the QGNN model at scale. Importantly, in this particular analysis, we employ a single QGCN circuit, so the absence of barren plateaus in the circuit also implies its absence in the QGNN framework. Here we conclude our analysis.
Figure 9: Average variance of the loss function partial derivatives $\overline{\operatorname{Var}}\left[\partial_{\theta_{\mu}}\mathcal{L}\right]$ of the QGCN circuit. The correlated QGNN (corr) uses the same parameters for all $W$ cells in a convolution layer, while in the uncorrelated (uncorr) all cells are independent. The variance for each variational parameter $\theta_{\mu}$ is calculated over 200 points in the loss landscape.
VII Conclusions
@@
We presented a novel Quantum Graph Neural Network framework that efficiently generates node embeddings by combining inductive representation learning with quantum circuit design. Specifically, our QGNN integrates sampling and aggregation techniques from GraphSAGE, using QGCNs as aggregators. The QGCN circuit is inspired by previous QCNN designs [lukin_qcnn_2019], featuring convolutional and pooling layers, while maintaining high flexibility to accommodate diverse architectures. By defining the convolutional cell, the circuit is uniquely determined, enabling the construction of tailored quantum circuits for specific problems or symmetry-preserving architectures.
@@
Benchmarking on the QM9 dataset shows that while the classical GNN achieves superior training performance, the QGNN exhibits stronger generalization, particularly as molecular complexity increases. In this evaluation, we employ a single, minimal QGCN circuit reused at every aggregation step. Hence, the quantum circuit architecture remains unchanged within each case, with tunable parameters staying constant regardless of the number of nodes in a graph, showcasing the ability to handle graphs with varying number of nodes without requiring design modifications in the circuit. Additionally, the quantum approach shows faster initial learning, indicating potential for rapid convergence. To further address trainability limitations observed in the single-circuit setup, we also investigate an extension of the proposed QGNN framework in which each aggregation step employs a distinct quantum circuit. In this configuration, different QGCN circuits are assigned to each hop, which remarkably boosts training performance across all tested cases.